OpenStudy (anonymous):

Is the answer to this problem, √3:(y^(3)) Y?

OpenStudy (anonymous):

yes it is y

OpenStudy (anonymous):

thought so. thank you.

OpenStudy (anonymous):

√3:(y^(3)) is the same thing as saying (y^3)^(1/3). Using the rule of powers, it simplifies to: y^(3 x 1/3) = y^1 = y. So that's correct.